An alpha particle has a mass of 6.62x10-27 kg and a typical speed of 1.6x10^7 m/s.
Show the calculation for the kinetic energy of the alpha particle in joules
1
Expert's answer
2019-11-05T07:40:22-0500

Answer: "8.47\u00d710^{-13}J"

Solution:

"E=\frac{m×V^2}{2}=\frac{6.62×10^{-27}kg×{({1.6×10^7}m/s)}^2}{2}=8.47×10^{-13}J"